OpenStudy (anonymous):

How long does it take your money to double at 8% interest?

OpenStudy (anonymous):

A close approximation is the "rule of 72" which is that the product of the interest rate and the time needed to double is 72 so 72 = 8 x and x = 9. Else, (1+0.08)^x = 2 take log of both sides x ln(1.08) = ln 2 x = ln(2)/ln(1.08) = 9.006

OpenStudy (anonymous):

they want me to use the formula... \[A=Pe^{rt}\]

OpenStudy (anonymous):

Doubling means that A'/A = 2 A'/A = exp(rt')/exp(rt) = 2 ln 2 = rt' - rt = r (t'-t) set t = 0 to start ln 2 = r t' = 0.08 t' t' = ln 2 / 0.08 = 8.66 The difference (9 vs. 8.66) is because the exp(rt) equation assumes continuous compounding and the (1+r)^t assumes compounding only at unit intervals of t, like years.

OpenStudy (anonymous):

Glad to have been of help.

OpenStudy (anonymous):

okay why does A= exponent rt?

OpenStudy (anonymous):

how would we plug into the formula... meaning what does each variable represent in this case in the formula I gave you.

OpenStudy (anonymous):

The amount = principal times growth factor let N be the number of dollars in the account dN/dt = N r growth equation for rate r dN/N = r dt then integrate from t=0 to t ln N - ln No = r t express as exponential N = No exp(rt) A = P exp(rt) in financial notation

OpenStudy (anonymous):

Okay , I don't think I 'll ever fully understand how to set this up, but thanks for the help.

OpenStudy (amistre64):

A is the amount you get after some time t ... A = Pe^rt; well P is doubled when A = 2P 2P = Pe^rt ... divide and log out